OpenStudy (anonymous):

PLEASE HELP AND WILL GIVE A MEDAL! Please need help!! Subtract and simplify 4a+1/a^2-4-3a+4/a^2-4 A. 8/a^2-4 B. a-3/a^2-4 C. 3/a+2 D. 2a-6/a^2-4

OpenStudy (solomonzelman):

I am guessing the question is, \(\normalsize\color{blue}{ \frac{ \LARGE{4a+1} }{\LARGE{a^2-4}} -\frac{ \LARGE{3a+4} }{\LARGE{a^2-4}} }\) right ?

OpenStudy (anonymous):

I thin so I can't really read what you put there my iPad is acting weird

OpenStudy (anonymous):

Oh it's showing up now. Yes it is

OpenStudy (solomonzelman):

4a+1 3a+4 ──── ─ ─── a²-4 a² - 4

OpenStudy (solomonzelman):

You have the same denominators, so you can add them.

OpenStudy (solomonzelman):

I mean subtract them.

jagr2713 (jagr2713):

is there an answer choice

jagr2713 (jagr2713):

its a pretty long explanation

OpenStudy (anonymous):

subtract in the numerator, leave the denominator alone

OpenStudy (anonymous):

\[4a+1-(3a+4)=4a+1-3a-4=a-3\]

OpenStudy (anonymous):

the denominator stays the same, you get \[\frac{a-3}{a^2-4}\]